OpenStudy (anonymous):

graph the equation and identify the y intercept x+5y=15 I have already graphed it the y intercept needs to be an ordered pair is it (0, 15)

jimthompson5910 (jim_thompson5910):

To find the y-intercept, simply plug in x = 0 and solve for y x+5y=15 0+5y=15 5y = 15 y = 15/5 y = 3 So when x = 0, y is y = 3 This gives us the y-intercept as the ordered pair (0, 3)

jimthompson5910 (jim_thompson5910):

You have your thinking backwards, but I think you have the right idea